OpenStudy (anonymous):

Solve for X and Y: 3x / 5 + y/2 =13/10 and .5y+ .04x = 2.42 then evaluate xy. a. xy=6 b. xy= -16 c. xy=30 d. xy= -10 e. none of these

OpenStudy (here_to_help15):

@Preetha help please

OpenStudy (here_to_help15):

@jigglypuff314

OpenStudy (anonymous):

@Here_to_Help15 Thank you <3 you're the best.

OpenStudy (here_to_help15):

anyone?? lol @cwrw238

OpenStudy (here_to_help15):

thanks ^^

OpenStudy (anonymous):

they dont seem to want to help :/

OpenStudy (here_to_help15):

@sammixboo

OpenStudy (here_to_help15):

@HelpBlahBlahBlah

OpenStudy (here_to_help15):

@skullpatrol

OpenStudy (here_to_help15):

come on _--_

OpenStudy (anonymous):

i can help

OpenStudy (here_to_help15):

then help her please

OpenStudy (here_to_help15):

-_-

jigglypuff314 (jigglypuff314):

got it multiply the first equation through by 10 multiply the second equation through by 100 you will get 6x + 5y = 13 50y + 4x = 242 follow so far?

OpenStudy (anonymous):

sorry, a guy with a similar name told me to not just give out answers

OpenStudy (anonymous):

i need to explain it seems

OpenStudy (anonymous):

yes so far im following jigglypuff

OpenStudy (anonymous):

so now you have to back substitute

OpenStudy (anonymous):

or just multiply the 2 equations so that the x or y has the same coefficient

jigglypuff314 (jigglypuff314):

cool :) then try doing elimination method... multiply the first equation through by 2 and the second by 3 and get 12x + 10y = 26 12x + 150y = 726

jigglypuff314 (jigglypuff314):

then we can subtract like that 12x - 12x = 0 150y - 10y = 140y 726 - 26 = 700 so 140y = 700 solve for y

jigglypuff314 (jigglypuff314):

then once you get y plug it back into an original equation... then solve for x

OpenStudy (anonymous):

okay let me do that and see what i get. hold on

OpenStudy (anonymous):

so y=5

OpenStudy (anonymous):

x=-2 ?

OpenStudy (anonymous):

so -10 ?

jigglypuff314 (jigglypuff314):

yes :) then xy = x*y = (-2)*(5) = -10 so correct! :)

OpenStudy (anonymous):

thank you so much!!

jigglypuff314 (jigglypuff314):

you're welcome :) glad I could help ^_^